OpenStudy (anonymous):

write a third degree polynomial function in standard form with 2 as a root

OpenStudy (lasttccasey):

x^3-8=0 Knowing when you plug 2 in, it will be cubed into an 8 so just subtract 8 to make it equal zero: x^3=0 x=2 so (2)^3 = 8 make it balance so bring the 8 to the left side for standard form. x^3-8=0 when x=2.

OpenStudy (anonymous):

thank you

OpenStudy (lasttccasey):

You're welcome.